OpenStudy (anonymous):

6th term of expansion (1+x)^32 quickest method

OpenStudy (campbell_st):

for a binomial \[(1+ x)^n = \sum_{r = 0}^{n}\left(\begin{matrix}n \\ r\end{matrix}\right) 1^{n - r}x^r\] so to find the 6th term let r = 5 remember the terms start at r = 0 0, 1, 2, 3, 4, 5 and evaluate \[\left(\begin{matrix}32 \\ 5\end{matrix}\right) 1^{32-5}\times x^5\]

OpenStudy (anonymous):

And what if the terms were (a+b) instead of (1+x)

OpenStudy (anonymous):

You were helpfull